A radioactive tracer used in a medical test has a half-life of 2.6 h. What proportion of this tracer
will remain after 24 h?
1
Expert's answer
2020-06-04T09:58:25-0400

The number of atoms "N" remaining at time "t" is given by the law od radioactive decay:


"N = N_02^{-\frac{t}{T}}"

where "N_0" is number of atoms at the initial moment of time, "T" is the half-life. Thus, the proportion of this tracer after 24 h will be:


"p = \dfrac{N}{N_0} = 2^{-\frac{t}{T}} = 2^{-\frac{24h}{2.6h}} = 2^{-9.23} = 0.0017"

Answer. p = 0.0017.